### v3

@@ 1 Introduction
While there is no general agreement among economists on whether a carbon tax or an ETS is better (Stavins,, 2022), the literature have often considered these two pricing instruments (trading versus taxes) as policy substitutes. For instance, some studies support carbon tax over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003). For instance, Newell and Pizer, (2003) find that carbon taxes can yield higher welfare benefits than an ETS under such uncertainties. Conversely, other studies advocate ETS over carbon tax because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009). Harstad and Eskeland, (2010), Hahn and Stavins, (2011), and Stavins, (2022) argue that from a practical perspective, ETS may be a preferred instrument over tax, as free allowances could be negotiated among participating agents to redistribute burden and, thus be used to gain political support. Knopf et al., (2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn,, 2008). In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016). However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms results in distortions in the market allocation. For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022). This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary.
2 Optimal Carbon Tax under the ETS
While carbon taxes and ETS are often considered as substitutable policies, we demonstrate that each country needs to implement an optimal carbon tax as a complementary policy within a global ETS regime to achieve optimal level of emissions. Before introducing the full multi-region dynamic general equilibrium model, it is helpful to first explore the optimal carbon tax within a static framework. Under the global ETS framework, we assume that each region $i$ is allocated with an exogenous level of emission cap ($\bar{E}_{i}$), which is the upper limit of emissions allowed for that region and is strictly enforced. The region’s gross emissions before abatement is $E_{i}^{\mathrm{Gross}}$, its emissions net of abatement is $E_{i}$, the amount of emission permits purchased (or sold) in the market is $E_{i}^{P}$ (if $E_{i}^{P}<0$, it implies that the region sells emission permits) so we have the emission cap constraint $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$, where the Nash equilibrium permit price is $m$ under the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$. Here we assume that each region is a price taker. The region’s marginal abatement cost (MAC) function, denoted $\mathrm{MAC}_{i}(E_{i})$, is assumed to be monotonically decreasing over emissions $E_{i}$ with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$. A social planner of the region solves the following minimization problem
@@ 2 Optimal Carbon Tax under the ETS
where the first term of the objective function is the regional damage from the global emissions $\sum_{j\in\mathcal{I}}E_{j}$ (which is the present value of future damage from the global emissions), the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E_{i}^{\mathrm{Gross}}-E_{i})$. For simplicity, we assume that emissions cannot be negative in our model.
In the static Nash equilibrium of all regional social planners, every region solves equation (1) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$. Let $E_{j}^{*}$ be the Nash equilibrium emissions for each region $j$ and let $m^{*}$ be the Nash equilibrium price. Thus, from the Karush-Kuhn-Tucker conditions of (1), the solution to the region’s optimal emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
$D_{i}^{\prime}\left(\sum_{j\in\mathcal{I}}E_{j}^{*}\right)=\mathrm{MAC}_{i}\left(E_{i}^{*}\right)-m^{*},$ (2)
when $E_{i}^{*}>0$. The marginal regional damage can also be regarded as the regional social cost of carbon (SCC). Note that in the static Nash equilibrium with $E_{i}^{*}>0$, the permit price $m^{*}$ should be equal to the shadow price of the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$. That is, if the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ is not binding, or $\sum_{j\in\mathcal{I}}E_{j}^{*}<\sum_{j\in\mathcal{I}}\overline{E}_{j}$ (i.e., there is an over-supply of emission permits), then the permit price is zero. Alternatively, if the permit price is nonzero, then the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ must be binding, so the supply of emission permits is equal to the demand (i.e., $\sum_{j\in\mathcal{I}}E_{j}^{*}=\sum_{j\in\mathcal{I}}\overline{E}_{j}$). We emphasize that the assumption commonly used in the literature—that the MAC equals the permit price (e.g., Keohane,, 2009; Nordhaus, 2010b, )—however, does not hold in this case, unless the regional marginal damage or SCC is zero, according to equation (2). Moreover, we also numerically verify in our dynamic Nash equilibrium that, when $E_{i}^{*}>0$, the regional MAC is also strictly larger than the permit price, and the difference is the regional SCC.
An important issue arises when a regional social planner cannot directly enforce its optimal emissions level on firms, the actual emission producers in the economy. To address this, let us assume that the social planner can employ carbon taxation as a policy instrument to resolve the discrepancy in optimal emissions between the firms and the regional social planner. Consider a representative firm in region $i$ that produces emissions and operates under the region’s MAC function and emissions cap constraint in a static framework. Since the firm does not internalize the damage from emissions, we assume that the social planner of region $i$ levies a tax on the firm of $\tau_{i}$ per unit of emissions. The firm solves the following minimization problem:
@@
where the first term of the objective function is the total tax on emissions. In the static Nash equilibrium of all representative firms, every firm solves equation (3) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$. Let $E_{j}^{*,\mathrm{Firm}}$ be the Nash equilibrium emissions for each firm $j$ and let $m^{*,\mathrm{Firm}}$ be the Nash equilibrium price. Thus, from the Karush-Kuhn-Tucker conditions of (3), the solution to the optimal emissions of the firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}},$ (4)
when $E_{i}^{*,\mathrm{Firm}}>0$. If the regional social planner imposes the tax to be equal to the regional SCC, i.e., the regional marginal damage, then from equations (2) and (4), we have $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ and $m^{*,\mathrm{Firm}}=m$. In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal emissions and the market price of emission permits, capturing the idea that the regional social planner can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its emissions level of $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$. When $E_{i}^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (3) shows that
$\tau_{i}\geq\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}}.$ (5)
That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits. In particular, the regional $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})$ at $E_{i}^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m^{*,\mathrm{Firm}}$, making the right hand side of the inequality (5) to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ if it is positive.
Therefore, for any regional social planner’s optimal emissions $E_{i}^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
$\tau_{i}=\max\{0,\ \mathrm{MAC}_{i}(E_{i}^{*})-m^{*}\},$ (6)
such that the firm’s optimal emissions are consistent with the social planner’s (i.e., $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$). The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model. Note that this tax is contingent on the imposition of the emission cap on the firm, and the tax level itself cannot make the firm’s emissions to be $E_{i}^{*}$ unless $m^{*}=0$.
It can be easily shown that our argument is consistent with the existing literature when there is no ETS. The cost minimization problem of the representative firm facing the emission cap regulation without the ETS becomes:
[…]
From the Karush-Kuhn-Tucker conditions of (7), when the firm’s optimal solution $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ is not binding at its lower or upper bound, we have the well known equality:
$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*}).$ (8)
This implies that the optimal carbon tax should be set equal to the MAC at the regional social planner’s optimal emissions $E_{i}^{*}$, provided $E_{i}^{*}$ is not binding. This result has been widely used in the literature (e.g., Nordhaus,, 2014; Cai et al.,, 2017; Cai and Lontzek,, 2019). Note that our formula (6) under an ETS with zero permit price is consistent with (8) under no ETS.
A regional social planner in each region $i$ has two targets: one is its exogenous NDC commitment $\overline{E}_{i}$, and the other is the regional social planner’s endogenous optimal emissions $E_{i}^{*}$ after permit trading (note that $\overline{E}_{i}$ and $E_{i}^{*}$ are unequal in general [^5] We believe this is a realistic consideration, as emission cap pathways may not always align precisely with regional optimal emissions over time. This discrepancy arises because the negotiation of emission permit allocations is often time-consuming and influenced by political considerations, which may not lead to economically optimal outcomes. Our model simulation shows that the emission cap pathways constructed from the Paris Agreement and the Glasgow Climate Pact indeed deviate from the optimal regional emissions.) in the Nash equilibrium of all 12 regional social planners. That is, the region $i$’s emissions should be $E_{i}^{*}$, which should also satisfy $E_{i}^{*}\leq\overline{E}_{i}+E_{i}^{P}$ after permit trading in the Nash equilibrium. Several policy instruments can be applied to enforce the firms to meet the targets, assuming that the firms can buy or sell emision permits. One such instrument is a quantity policy, in which the regional social planner distributes $\overline{E}_{i}$ (e.g., by means of an auction or grandfathering) to its firms, ensuring the firms’ aggregate emissions, $E_{i}^{*,\mathrm{Firm}}$, do not exceed the exogenous allocated cap after permit trading (i.e., $E_{i}^{*,\mathrm{Firm}}\leq\overline{E}_{i}+E_{i}^{P}$) in the Nash equilibrium of firms of all regions. [^6] The representative firm framework can be generalized to an environment with multiple firms. Assume that there are $n$ homogenous firms in the region $i$, each having $\overline{E}_{i}/n$ as their caps, with a firm’s MAC function at its emissions $e$ is defined as $\mathrm{MAC}_{i}(ne)$, where $\mathrm{MAC}_{i}(\cdot)$ is the aggregate regional MAC function used by the regional social planner and $E=ne$ is the aggregate regional emissions. Under this definition, the aggregate total abatement cost of all firms in region $i$ is equal to the integral of $\mathrm{MAC}_{i}(E)$ on $[E_{i},E_{i}^{\mathrm{Gross}}]$, when each firm’s emission is $E_{i}/n$ while $E_{i}^{\mathrm{Gross}}/n$ is each firm’s upper bound under no abatement. In this case, the Nash equilibrium of the economy with multiple firms is identical to the equilibrium with a representative firm. However, under a quantity-only control policy, $E_{i}^{*,\mathrm{Firm}}$ is not equal to $E_{i}^{*}$, as the zero-tax Nash equilibrium of the firms, $\left\{m^{*,\mathrm{Firm}},E_{i}^{*,\mathrm{Firm}}:j\in\mathcal{I}\right\}$, is not the Nash equilibirum of the regional social planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$. This becomes particularly evident in the following extreme case: assume that $\overline{E}_{i}$ is larger than $E_{i}^{\mathrm{Gross}}$ for every region, which leads to $m^{*}=0$ under the ETS. The firms’ Nash equilibrium without a tax yields $E_{i}^{*,\mathrm{Firm}}=E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})=m^{*,\mathrm{Firm}}=0$. In contrast, the regional social planners’ Nash equilibrium leads to $E_{i}^{*}<E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*})>m^{*}=0$ in general, similar to a model without emission caps. Note that this quantity policy cannot directly distribute $E_{i}^{*}$, as doing so would either result in the region violating its commitment if $E_{i}^{*}>\overline{E}_{i}$ or losing the benefit in selling permits if $E_{i}^{*}<\overline{E}_{i}$. Another instrument is a carbon tax policy, in which the region $i$ imposes a carbon tax $\tau_{i}$ on the firms, e.g., $\tau_{i}=\mathrm{MAC}_{i}(\overline{E}_{i})$ or $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})$. However, a carbon tax policy alone can target only one emission level, either $\overline{E}_{i}$ or $E_{i}^{*}$, in each region, while $\overline{E}_{i}\neq E_{i}^{*}$ in general. Therefore, neither instrument can achieve both targets simultaneously.
We argue that in the presence of a global ETS, it is optimal for each region $i$ to impose both the quantity policy and an additional carbon tax policy simultaneously on firms to enforce both targets. Here the additional carbon tax for region $i$ is chosen to be $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ when $E_{i}^{*}>0$, under the Nash equilibirum of the planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$. If some regions impose only the quantity policy while the others impose both policies, this leads to a Nash equilibrium of firms, which is not the Nash equilibirum of the planners. Note that a tax has no impact on the regional social planners’ decisions; $E_{i}^{*}$ remains the optimal decisions of the regional social planners regardless of the tax levels. However, the tax can influence the firms’ decisions. Therefore, from the regional social planners’ perspective, any Nash equilibrium of firms deviating from $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$ is worse, and the regions should always impose both the quantity policy and the optimal additional carbon taxes simultaneously on the firms to meet both targets. [^7] Our argument is based on the assumption that the emission caps are exogenous in the regional level.

### v4

@@ 2 Related Literature
While there is no general agreement among economists on whether carbon taxation or ETS is better (Stavins,, 2022), previous works have often considered these two pricing instruments (trading versus taxes) as policy substitutes. For instance, some studies support carbon taxation over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003). For instance, Newell and Pizer, (2003) find that carbon taxation can yield higher welfare benefits than ETS under such uncertainties. Conversely, other studies advocate ETS over carbon taxation because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009). Harstad and Eskeland, (2010), Hahn and Stavins, (2011), and Stavins, (2022) argue that from a practical perspective, ETS may be a preferred instrument over tax, as free allowances could be negotiated among participating agents to redistribute burden and, thus be used to gain political support. Moreover, in a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016). However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms may result in distortions in the market allocation. [^5] Environmental pollution has also been studied from the context of fiscal federalism, which considers what level of government should regulate pollution (Oates,, 1999; Ogawa and Wildasin,, 2009; Banzhaf and Chupp,, 2012; Williams III,, 2012). For a comprehensive comparison on carbon taxation and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022). This study contributes to the literature by showing that regional carbon tax is complementary to ETS under noncooperation.
3 A Static Framework of Climate and the Economy with an ETS
This section presents a simplified static model framework with a global ETS to provide intuition on the relationship between the regional SCC and regional MAC. In a global ETS regime, the main forces determining the equilibrium are climate damages, costs (or revenue) from emissions trading, and costs from regional emission abatement efforts. In this simplified static framework, it can be easily shown that the regional SCC equals the regional MAC minus the market price of emission permits. Building on these insights, the next section presents the multi-region dynamic general equilibrium model, followed by the quantification strategies and numerical analysis.
Consider a world economy with multiple regions under a global ETS regime, and let $\mathcal{I}$ denote the set of regions. Each region $i\in\mathcal{I}$ is allocated an exogenous emission cap $\bar{E}_{i}$, which represents the region’s maximum allowable emissions and is strictly enforced. Economic activity generates emissions, with $E_{i}^{\mathrm{Gross}}$ denoting a region’s gross emissions, which are treated as exogenous in this simplified framework. [^6] It is innocuous to assume that gross emissions are exogenous in this static framework. In the full dynamic model introduced later, regional gross emissions are proportional to GDP, which is determined by capital and the exogenous population growth. Given that the current level of capital is determined in the previous period, gross emissions can be considered exogenous for each period in a static equilibrium. To comply with the emission cap constraint, each region’s social planner can either (i) purchase or sell emission permits in the global emission market, or (ii) undertake costly abatement efforts. Let $E_{i}$ denote its emissions net of abatement (henceforth referred to as net emissions), and $E_{i}^{P}$ the quantity of emission permits purchased in the market. A positive $E_{i}^{P}$ indicates that the region purchases permits, while a negative $E_{i}^{P}$ indicates that the region sells permits. The emission cap constraint for each region is then given by $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$.
In this economy, the regional social planner aims to minimize the total economic costs associated with its emissions, including (i) climate damages due to global emissions, (ii) costs (or revenue) from emission trading, (iii) and costs from own regional emission abatement. Formally, a social planner of each region solves the following minimization problem:
@@ 3 A Static Framework of Climate and the Economy with an ETS
The first term of the objective function represents regional climate damages, where $D_{i}(\cdot)$ is a function that captures the regional climate damages from global net emissions, $\sum_{j\in\mathcal{I}}E_{j}$. Here $D_{i}(\cdot)$ can also be considered as the present value of future climate damages from the current global net emissions in the corresponding dynamic framework. The second term accounts for the cost (or revenue) from purchasing (or selling) permits at price $m$, which is the Nash equilibrium price satisfying the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$. The quantity of emission permits purchased may influence the global permit price, making the market price depend on their emission permit purchase choices, which in turn are contingent on the trade decisions of other regions. Lastly, the third term reflects the total abatement cost incurred to reduce emissions by $(E_{i}^{\mathrm{Gross}}-E_{i})$, where $\mathrm{MAC}_{i}(\cdot)$ denotes the region’s marginal abatement cost and is assumed to be monotonically increasing over $E_{i}$, with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$. For simplicity, it is assumed that net emissions are nonnegative throughout the paper. [^7] This assumption ensures that no region can abate more than its gross emissions solely to sell permits to other regions.
In the static Nash equilibrium of the world economy, every regional social planner simultaneously solves equation (1) while satisfying the market-clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies that $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$. We assume that each region is a price taker for permits, as a large number of firms in each region participate in the permit trading market in practice. Thus, the Karush-Kuhn-Tucker conditions of equation (1) imply that the solution to the region’s optimal net emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
@@
when $E_{i}>0$. The left hand side of this equation represents the marginal regional damage, which corresponds to the regional SCC. This term captures the additional regional economic cost imposed by a unit increase in global emissions, reflecting the region’s contribution to global climate change. The right hand side captures the difference between $\mathrm{MAC}_{i}(E_{i})$ and $m$, reflecting the deviation of regional abatement efforts from the market equilibrium. The equation (2) simplifies to $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$, where $\mathrm{SCC}_{i}$ denotes the regional SCC. A positive regional SCC implies $\mathrm{MAC}_{i}(E_{i})>m$, suggesting that a region experiencing climate damages has an incentive to abate emissions to a level where the marginal abatement cost exceeds the market price of emission permits.
In the environmental economics literature, it is well established that when production generates emissions as a negative externality in an economy, a regulator can achieve the socially efficient level of abatement by setting an emission cap at that level and issuing tradable permits to firms. The resulting market equilibrium under cap-and-trade is equivalent to the outcome of an optimal emission tax equal to the MAC at the efficient abatement level, in the absence of cost uncertainty or transaction costs. However, in a multi-region framework where climate damages are region-specific, the decentralized equilibrium outcome under the global ETS does not necessarily coincide with the regional social planner’s optimal outcome. The underlying reason is that, because climate damages are heterogeneous across regions, a uniform global carbon price alone may not yield regionally optimal levels of emission abatement or the corresponding emission prices.
The intuition behind $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ can be further illustrated through an extreme case. If emission caps are sufficiently large such that they are not binding for any region—effectively a scenario without emission constraints—permit prices and trading volumes would be zero. Imposing the equilibrium condition $\mathrm{MAC}=m$ in this situation would imply no abatement (i.e., $\mathrm{MAC}=0$ when $m=0$). However, it is well established that even in the absence of binding caps, the optimal level of emission abatement from the perspective of a regional social planner remains strictly positive (i.e., $\mathrm{SCC}=\mathrm{MAC}>0=m$) under noncooperative Nash equilibrium (see, e.g., Nordhaus and Yang,, 1996; Cai et al.,, 2023). This example further highlights the gap between the regional social planner’s optimal abatement level and the decentralized market outcome, as firms undertake no abatement in the absence of an emission price. The basic intuition explored here carries over to the dynamic general equilibrium model introduced in the next section, where we further confirm numerically that $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ continues to hold.
An additional important implication of this relationship is that the two pricing instruments—emission taxes and cap-and-trade—are not competing policies but rather complementary under the global ETS framework. Although our model does not explicitly introduce firms, the regional MAC is effectively obtained as an aggregation of individual firms’ marginal abatement quantities at a given price level within a region (Keohane and Olmstead,, 2016). To achieve the optimal level of net emissions, $E_{i}$, from each regional social planners perspective under the global ETS, the optimal policy for each regional planner is to impose, within its own economy, a carbon tax equal to its regional $\mathrm{SCC}_{i}$ on top of the permit price from the global ETS. This ensures that individual firms internalize both the regional SCC and the global permit price. With the regional tax incorporated, each individual firm chooses net emissions such that its own MAC equals the sum of the permit price and the regional carbon tax, i.e., $\mathrm{MAC_{i}}(E_{i}^{\prime})=\mathrm{\tau}_{i}+m^{\prime}$, where the optimal carbon tax is $\mathrm{\tau}_{i}=\mathrm{SCC}_{i}$ from the perspective of the regional social planner.
For a partial ETS, if region $i$ participates in the ETS, then the relation $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ still holds, where $m$ is the equilibrium permit price between the participating regions; if region $i$ does not participate, then the relation is changed to $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})$, which is a standard relation under no ETS.